OpenStudy (anonymous):

A golf ball is hit from the ground with an initial velocity of 208 feet per second. Assume the starting height of the ball is 0 feet. How long will it take the golf ball to hit the ground?

OpenStudy (anonymous):

Do you have any initial trajectory information (i.e. angle at which the ball is hit)?

OpenStudy (anonymous):

13secs

OpenStudy (anonymous):

h(t) = vo t - 1/2 gt^2 where vo = 208 ft/s and g = 32ft/s^2 h(t) = 208 t - 16t^2 when ball hits the ground h(t) = 0 208 t - 16t^2 = 0 => 16 t ( 13 - t) = 0 t = 0 and 13 sec 0 s is the time when ball is hit from the ground ball hits the ground after 13 s

OpenStudy (amistre64):

your vo t is off

OpenStudy (amistre64):

vo sin(a) t would be the upward force involved in this so there are at most 2 answers

OpenStudy (amistre64):

or maybe infinitely many ...

OpenStudy (amistre64):

we need either a distance; or some other peice of information